# OnlyFlow: Optical Flow based Motion Conditioning for Video Diffusion Models

### v1

@@ Abstract
color=yellow!60,inline]H:regle du pouce, ajouter texte entre section et sous sections Dire dans les RW, si methode comparable, expliquer puis critiquer, et pourquoi notre modele a lieu d’être)
We consider the problem of text-to-video generation tasks with precise control for various applications such as camera movement control and video-to-video editing. Most methods tacking this problem rely on providing user-defined controls, such as binary masks or camera movement embeddings. In our approach we propose OnlyFlow, an approach leveraging the optical flow firstly extracted from an input video to condition the motion of generated videos. Using a text prompt and an input video, OnlyFlow allows the user to generate videos that respect the motion of the input video as well as the text prompt. This is implemented through an optical flow estimation model applied on the input video, which is then fed to a trainable optical flow encoder. The output feature maps are then injected into the text-to-video backbone model. We perform quantitative, qualitative and user preference studies to show that OnlyFlow positively compares to state-of-the-art methods on a wide range of tasks, even though OnlyFlow was not specifically trained for such tasks. OnlyFlow thus constitutes a versatile, lightweight yet efficient method for controlling motion in text-to-video generation.
Models and code will be made available on GitHub and HuggingFace.
Figure 1: OnlyFlow controls the generation of video with text and the motion of a video input, using an estimation of its optical flow. We highly encourage readers to check our supplementary content for video results, which cannot be well represented by still images.
color=teal!60,inline]M: alpha 0.5 * 2 flow superposition
CVPR 2023 transfer optical flow: https://diffusion-motion-transfer.github.io/MotionEditing_arXiv.pdf
fig 2 : changer RAFT
loss de training dans related work et citer dans partie 4
alpha en parametre dans la fig 2
3 Framework / architecture la plus generale possible 3.1 pas citer directement RAFT Plus à rework dans partie training 4 Training 5 Experimental results
metric intro dans fin de partie 4
@@ 1 Introduction
There has been tremendous progress in generative AI thanks to the rise of diffusion models for text-to-image (T2I) with StableDiffusion [10]. More recently T2I have been extended to text-to-video (T2V) models (such as AnimateDiff [14]). This has started revolutionizing the creation industry (entertainment professional, advertising, art) by offering increased productivity and creativity. Text-based models do not offer precise user control, as the text modality often limit the user expression. Image-to-video models [20, 15, 7] help to condition the start of the video but do not provide long range temporal conditioning. Many recent work aims to provide more precise control with the use of external conditioning data for T2I [43, 25, 41] and T2V models [23, 9, 37, 42, 13]. Users can edit parts of the generated content, with the use of masks, have style customization from existing images or even have control between the conditioning of content, layout or style.
@@
Figure 2: Overview of OnlyFlow. Inputs are i) a tokenized and encoded text prompt, ii) noisy latents for the diffusion model and iii) the optical flow of an input video. The latter is fed through a trainable optical flow encoder which outputs features maps that are injected in the diffusion U-Net. We experiment with several injection strategies, for illustration purposes we only show the injection in temporal attention layers of the U-Net. The U-Net is kept frozen during training. The output generated video matches the prompt and the auxiliary video’s motion.
The creation of videos presents a specific challenge linked to the complexity of motion synthesis. By integrating additional motion control, one can reduce the ambiguities inherent in video synthesis. This allows for better model movements and improved ability to manipulate generated content for personalized creations.
In this paper, we introduce OnlyFlow, a simple solution to control video generation from motion. This consists in passing an auxiliary video as input to copy its motion. A simple way to describe the motion in a video is the optical flow. This description of the temporal evolution of pixels is widely used in different tasks in computer vision. OnlyFlow extracts optical flow from a given video and employs it to guide the denoising diffusion process of a video generation model based on AnimateDiff using a small external trained optical flow encoder.
@@
To summarize, (1) we propose OnlyFlow, a new motion-guided video generation model that enables motion conditioning from given reference videos; (2) we design a motion control strategy based on optical flow representation to perform motion guidance, allowing for motion transfer across videos; (3) we present metrics on its text alignment, motion fidelity, and temporal consistency. (4) we empirically validate the effectiveness and versatility of OnlyFlow in various video generation tasks, in which several experiments demonstrate its use case.
color=teal!60,inline]M: tres tres general au debut comme l’avancée des modèles de text to image et text to video + applications et sujets d’application
questionnement de controles très important et les modeles out of the box ne permettent pas ca ainsi beaucoup de modeles ont ete apportés … citer pour donner du contrôle (ce qui s’applique aux 2 type images et videos) enumerer les types de controle - edition des parties depuis un masque - prolongement de video - influencer avec un style via image (IP-adapter) - ou composition ou layout contenu
On s’interesse aux modèles de video dont pleins de méthodes on été développées pour conditionner par d’autres inputs generalement (revenir aux autres approches), comment ? Structure commune avec un modele annexe (copie pour cn, encodeur pour t2i, etc..)
user-defined conditioning (controlnet + vectorial representation of camera movement)
mais …. pbs user defined .. dev les pbs pas facilement tedious task et ne permet pas des choses simples (ex conditionné par le mouvement, ce qui nous interresse ici)
ainsi l’objet qui nous permet d’isoler le mouvement d’une vidéo est le flow (phrase bateau sur l’utilisation du flow, maintes utilisations)
on a pris inspiration des techniques qui prenenent un modele annexe et du flow pour introduire OnlyFlow.
We introduce…(à copier) ce qu’est le modèle comment il est train ses applications (text-to-video, video-to-video style transfer, personalizing..)
les contributiosns en bulletpoint - Oui, utilisation du flow pour conditiioner animatediff - ses applications Montrer framework unique pour différentes types de tâches de customization: - ok style transfer - ok pour type contrôlnet - controle de camera - Modèle en open source, qui s’integrère facilement des plugins existant de par sa nature
color=yellow!60,inline]H:Truc a garder que j’ai tej de l’intro :
A common way to inject conditioning is to make use of an external trainable model tailored for a specific task. This model can be of the same size as the main T2I/T2V model, like in the case of the ControlNet [43], or much smaller in the case of IP-Adapter [41] and T2I-Adapter [25]
While existing models like VideoPoet [24] utilize optical flow for motion control, their weights are not publicly accessible.
@@ 2.1 Text-to-video diffusion models
Recent advances in text-to-image (T2I) generation have been propelled by sophisticated text encoders, enabling high-quality image synthesis from textual descriptions. Building on these successes, there’s growing interest in text-to-video (T2V) generation, aiming to extend these capabilities to generate coherent video content.
To leverage powerful pre-trained image generators like Stable Diffusion [30], some approaches introduced temporal layers [14] into existing 2D architectures and fine-tuned these models on large video datasets.
VideoLDM [4, 18] introduced a motion module that uses 3D convolutions and temporal attention to capture frame-to-frame correlations effectively. AnimateDiff [14] enhanced a pre-trained T2I diffusion model by fine-tuning specialized temporal attention layers on video data, adding motion capabilities without altering the original image generation process. This approach allows for high-quality animation creation while retaining the strengths of the base T2I model. Moreover, animatediff is compatible with all existing T2I tools such as IP-Adapter [41] and Controlnet [43].
Other methods like Align-Your-Latents [5] convert T2I models into video generators by aligning noise maps across frames. Lumiere [3] improves temporal coherence by generating videos directly at the desired frame rate, bypassing temporal super-resolution modules. Techniques such as W.A.L.T [16]. and Sora [6] operate in compressed spatiotemporal latent spaces to handle video data efficiently. Addressing data scarcity, Videocrafter2 [8] learns motion patterns from low-quality videos while capturing appearance details from high-quality images. Nevertheless, these methods lack a mature ecosystem of control and edition tools, which in practice is the reason for their slow adoption.
@@ 2.2 Controllable video generation
Recent developments in controllable image generation have sparked interest in bringing similar control mechanisms to video generation. Researchers have introduced various control signals to guide the creation of videos, such as specifying the initial frame [15, 37], defining motion trajectories [37, 31, 42], focusing on particular motion regions, and controlling specific moving objects [38]. Some methods also use images as control inputs to enhance video quality and improve temporal relationship modeling [37, 24].
Another approach involves reference-based video generation, where motion from an existing real video directs the creation of new content. Techniques that integrate frame-wise depth or edge maps aim to regulate motion in a manner similar to control mechanisms used in text-to-image models [23, 43, 9, 22]. However, these methods can unintentionally introduce static features unrelated to the intended motion, leading to inconsistencies between the generated video and the desired content.
To overcome these challenges, motion-specific fine-tuning frameworks have been developed to extract distinct motion patterns from single videos or collections with similar motion [14, 39, 13]. While these methods show promise in capturing nuanced motion, they often involve complex training processes [39] or costly denoising (DDIM) process inversion [32, 23, 13] and can risk degrading the model’s performance.
Controlling camera movement during video generation is another area of focus. Some techniques involve fine-tuning models for different shot types using methods like MotionLoRA finetuning in AnimateDiff [14]. Others introduce camera representation embeddings to manipulate camera poses [17, 38], but these may be limited by the number of camera parameters they can handle or may require fine-tuning parts of the video diffusion model, which can affect the model’s generalization ability.
Overall, while significant progress has been made in controllable video generation, providing precise control over motion of intrinsic object and camera dynamics without compromising model performance remains an ongoing challenge.
color=yellow!60,inline]H: SD + LDM / LVM
AnimateDiff
T2I-Adapter
User defined text-to-video conditioning ControlNet (c’est bien mais trop lourd) + CameraCtrl + MotionCrl
Papiers recents
@@ 3.1 Diffusion model for video generation
Text-to-video generation involves creating coherent and high-quality videos conditioned on textual prompts. Due to the high dimensionality and temporal complexity of video data, recent approaches project videos into a lower-dimensional latent space using an encoder $\mathcal{E}$ [30]. Given a batch $B$ of video $\mathbf{V}\in\mathbb{R}^{F\times H\times W\times 3}$ with $F$ frames, the latent representation is obtained as $\mathbf{z}_{0}=\mathcal{E}\left(\mathbf{V}\right)\in\mathbb{R}^{F\times\frac{H}{l}\times\frac{W}{l}\times C}$, spatialy reduce by a factor $l$. A decoder $\mathcal{D}$ reconstructs the video from the latent space, yielding $\hat{\mathbf{V}}=\mathcal{D}(\mathbf{z}_{0})$.
@@
To effectively capture spatial and temporal dynamics, $\epsilon_{\theta}$ is typically implemented as a 3D U-Net architecture augmented with temporal convolutions and attention mechanisms. Each feature map represents a 5D tensor of $B\times F\times H\times W\times C$. This 3D U-net is factorized over space and time, meaning that each layer only operates on the space or time dimension, leveraging spatial convolutions for frame-wise feature extraction and temporal convolutions to model motion across frames:
  • spatial: each old 2D convolution layer as in the 2D U-net is extended to be space-only 3D convolution over $H\times W\times C$. Each spatial attention block remains as attention over space
  • time: a temporal attention block is added after each spatial attention block. It performs attention over $F$. The temporal attention block is important to capture good temporal coherence.
@@ 3.2 Optical flow conditioning in T2V Architecture
In this paper, we propose to modify this traditional T2V scheme by adding motion constraint. Fig. 2 illustrates an overview of our model architecture. More precisely, for enhanced controllability in video generation, we propose to condition the video generation on an optical flow signal, extracted from an auxiliary video provided to the system. We use a standard optical flow extraction $\mathcal{T}$ technique to obtain the optical flow $\mathbf{F}=\mathcal{T}(\mathbf{V})\in\mathbb{R}^{F\times H\times W\times 2}$ of the auxiliary input video, describing pixel displacements between frames, across $F$ frames.
The Architecture of OnlyFlow extends the denoising U-Net structure of a latent diffusion model by incorporating an optical flow encoder that processes optical flow input and extracts motion features at various scales. We then proceed to inject this signal into the attention layers of the U-Net. Let’s describe these two steps:
  • Motion feature extraction: the optical flow encoder $\Phi$ extracts motion features at different scales to incorporate conditioning signals in the denoising U-Net. It initially unshuffles the optical flow to increase the number of dimensions before applying a convolutional layer. Subsequently, for each block of the U-Net, it outputs optical flow representation embeddings using a series of ResNet blocks and temporal attention mechanisms. In the temporal attention layer, incoming tensors of shape $B\times F\times H\times W\times C$ are reshaped to $(B\times H\times W)\times F\times C$ before performing attention operations across the temporal axis. This is architecture similar to that used in T2I-adapter [25] and CameraCtrl [17].
  • Motion feature injection: with the multi-scale optical flow features extracted, we inject them into the U-Net of the T2V model. The whole process is summarized in Fig. 3: we combine the latent features $h_{k}$ of the k-th block of the U-Net with the control signal from flow features $C_{k}$ through element-wise addition. The combined feature is then processed by a linear layer and scaled using a parameter $\gamma$, and its output is fed directly into the first temporal attention layer of each motion module:
    ${h_{out}}_{k}=h_{k}+\gamma\times{W_{\theta_{1}}}_{k}\left(C_{k}+h_{k}\right).$ (2)
The parameter $\gamma$ effectively controls the influence of the optical flow from the input auxiliary video onto the generated video. It is the main parameter to improve or decrease the auxiliary motion injection over the generated output. We call $\gamma$ the optical flow conditioning strength.
@@
4 Implementation and training details
We describe the experimental protocol used to implement and train OnlyFlow.
4.1 Dataset and Preprocessing
OnlyFlow is trained on the WebVid dataset [1], which comprises of 10.7M video-caption pairs, totaling 52K hours of video content. The majority of these videos have a length between a few seconds and 1min, with 100k to 500k pixels by frame. All videos are spatially resized to $256\times 384$ pixels, and $F=16$ consecutive frames are randomly chosen, matching the training practice used in AnimatedDiff [14].
4.2 Model training and inference
Diffusion loss and training pipeline
Following the original formulation of video diffusion model training, the loss function is adjusted to include our control signal, i.e the optical flow $\mathbf{F}=\mathcal{T}(\mathbf{V})\in\mathbb{R}^{F\times H\times W\times 2}$ of the auxiliary input video $\mathbf{V}$. The optimization objective minimizes the mean squared error between the predicted noise and the actual noise:
$\mathcal{L}\left(\theta_{1},\theta_{2}\right)=\mathbb{E}_{\mathbf{z}_{0},\epsilon,c,t}\left[\left|\epsilon-\epsilon_{\theta_{1}}(\mathbf{z}_{t},p,\Phi_{\theta_{2}}(\mathbf{F}),t)\right|_{2}^{2}\right]$ (3)
where $\epsilon_{\theta_{1}}$ is conditioned on $c=\{p,\mathbf{F}\}$ containing the text prompt $p$ and the optical flow $\mathbf{F}$ with the use of a jointly trained encoder $\Phi_{\theta_{2}}.$
During the training phase, we freeze the image generation U-Net and the previously added temporal layers of AnimateDiff. Only the optical flow encoder and the merging linear layers in the attention blocks are trained, using an optical flow conditioning strength $\gamma=1.0$. To mitigate the distribution shift between the WebVid dataset and the dataset used to train the base T2I U-Net, we incorporate the domain adapter used in AnimateDiff training. This adapter is implemented as a LoRA [21] applied to the U-Net spatial layers.
Optical Flow extraction
We employed the commonly used RAFT-Large [33] model to generate a flow field $\mathbf{F}$, with the default values in RAFT inference settings. RAFT is a classic optical flow estimation model, it is fast during inference and widely used for computer vision tasks.
The base T2I model used in this work is the common Stable Diffusion 1.5 model composed of 4 block in its U-Net encoding and decoding part. Consequently, we use an encoder designed to output a corresponding number of optical flow conditioning signals $\Phi\left(\mathbf{F}\right)=\{C_{1},C_{2},C_{3},C_{4}\}$.
We opt to always insert the optical flow conditioning representations into the temporal attention blocks. This decision is motivated by the temporal attention layer’s ability to capture temporal relationships, which aligns with the sequential and causal nature of optical flow.
Spatial and temporal consistency
Nevertheless, it also makes sense to inject the features into the spatial attention layers. The spatial attention layers of the U-Net might adapt the generated visuals to the distortion brought by the optical flow conditioning. Thus, we also study a joint injection of features in both spatial and temporal attention layers, and we label this variation OnlyFlow-ST, while the variation with only temporal attention is labeled OnlyFlow-T.

### v2

@@ Abstract
We consider the problem of text-to-video generation tasks with precise control for various applications such as camera movement control and video-to-video editing. Most methods tackling this problem rely on providing user-defined controls, such as binary masks or camera movement embeddings. In our approach we propose OnlyFlow, an approach leveraging the optical flow firstly extracted from an input video to condition the motion of generated videos. Using a text prompt and an input video, OnlyFlow allows the user to generate videos that respect the motion of the input video as well as the text prompt. This is implemented through an optical flow estimation model applied on the input video, which is then fed to a trainable optical flow encoder. The output feature maps are then injected into the text-to-video backbone model. We perform quantitative, qualitative and user preference studies to show that OnlyFlow positively compares to state-of-the-art methods on a wide range of tasks, even though OnlyFlow was not specifically trained for such tasks. OnlyFlow thus constitutes a versatile, lightweight yet efficient method for controlling motion in text-to-video generation.
Figure 1: OnlyFlow controls the generation of video with text and motion of a video input, synthetically generated or not. We strongly encourage readers to check our supplemental content for video results that are not well represented by still images.
@@ 1 Introduction
Progress in generative AI has made tremendous progress thanks to the rise of diffusion models [20] and colossal datasets [35]. Previously, research on generative models focused mainly on generating images without any direct control, but today’s generative models have shifted towards integrating text to control the generation process and to facilitate interaction with the user. This has started to revolutionize the creation industry (e.g., professional entertainment, advertising, art) by offering increased productivity and creativity. However, today’s synthesis pipelines follow the text-based generative paradigm, which does not provide precise user control because the text modality often limits user expression.
In text-to-video (T2V) synthesis, the controllability problem is more pronounced than text-to-image (T2I), since the generation process has an additional axis: the temporal dimension. Yet, the generative community has made some progress in adding more control to T2V pipelines. Recent work has focused on providing more precise control by using external conditioning data for T2V models [23, 10, 42, 51, 13]. For example, users can edit parts of the generated content using masks, apply styles from existing images, or even switch between content, layout, and style conditioning. Despite that, the underlying challenge in video creation is related to the complexity of motion synthesis. By adding motion constraints, one can reduce the ambiguity inherent in video synthesis. This allows for better motion modeling and improved ability to manipulate generated content for personalized creations.
@@
Figure 2: Overview of OnlyFlow. Inputs are i) a tokenized and encoded text prompt, ii) noisy latents for the diffusion model and iii) the optical flow of an input video. The latter is fed through a trainable optical flow encoder which outputs features maps that are injected in the diffusion U-net. We experiment with several injection strategies, for illustration purposes we only show the injection in temporal attention layers of the U-net. The U-net is kept frozen during training. The output generated video matches the input prompt and motion.
In this paper, we introduce OnlyFlow, a simple solution for controlling video generation by directly integrating motion cues. In a nutshell, OnlyFlow constrains the generated video to mimic the input motion. More specifically, our proposed approach exploits optical flow information and uses it to guide the generation process using a small externally trained optical flow encoder. However, unlike most methods that use optical flow, our method works together as a T2V or video-to-video (V2V) model, as seen in Figure 1. On the one hand, OnlyFlow can use synthetically generated flows to simulate any movement, thus fixing the content to the text while respecting the input motion. On the other hand, we can use the flow of an input video and translate it to the generated one.
@@
To summarize, our contributions are threefold. (1) We propose OnlyFlow, a simple and novel motion-guided strategy for video synthesis that enables motion conditioning from given reference videos based on optical flow representations, allowing motion transfer across videos. (2) We conduct extensive empirical studies to validate OnlyFlow from both quantitative and qualitative perspectives. (3) We show the versatility of our approach in various video generation tasks, where several experiments demonstrate its use case.
@@ 2.1 Text-to-video diffusion models
Recently, T2V synthesis has been used to create realistic videos for creative applications. These applications build on generative backbones such as AnimateDiff [15], CogVideoX [22], or VideoCrafter [8]. A common technique for training video synthesis models is to use powerful pre-trained image generators [34]. To do this, most approaches have introduced temporal layers [15, 4, 19, 8, 37, 3] in the form of temporal attention layers and 3D convolutions, and fine-tuned these models on large video datasets [1, 43]. Other approaches [5, 46, 24] operate directly on the features of the T2I model. Finally, some methods train their method from scratch [14, 6, 22, 21, 17, 48]. These works typically tune their model in two steps: using a large amount of low quality data (videos and images) and then fine-tuning the model with high quality but scarce videos. However, all these models have a common flaw: motion controllability is limited. In this paper, we improve this aspect and build OnlyFlow on top of the popular AnimateDiff [15] model, but it can be easily extended to any T2V.
@@ 2.2 Controllable video generation
Recent developments in controllable image generation have sparked interest in bringing similar control mechanisms to video generation. Researchers have introduced various control signals to guide the creation of videos, such as specifying the initial frame [16, 42], defining motion trajectories [42, 36, 51], focusing on particular motion regions, controlling specific moving objects [44, 45], or introducing images as regularizers to enhance video quality and improve temporal relationship modeling [42, 26]. Additionally, many works propose motion-specific fine-tuning approaches [15, 46, 13]. While showing promise in capturing nuanced motion, they involve demanding training processes [46, 15] or costly inversion processes [38, 23, 13], risking degrading the model’s performance, while not being user-friendly. Finally, controlling camera motion during video generation is another area of focus. [18, 44] introduce camera representation embeddings to manipulate camera poses. But, these may be limited by the number of camera parameters they can handle. While camera control seems promising in theory, it is limited in terms of actual creation, as creating any complex motion is difficult.
Overall, while significant progress has been made in controllable video generation, providing precise control over motion of intrinsic object and camera dynamics without compromising model performance remains an ongoing challenge. We believe that OnlyFlow opens a direction towards achieving this goal.
2.3 Optical Flow as Motion Priors for Video Synthesis
Closest to our work are the studies that integrate optical flow into the generation pipelines. Undoubtedly, optical flow is an intuitive signal to represent motion in videos. Therefore, many works have explored the introduction of optical flow to guide the generation towards a specific motion, especially in V2V and I2V generation [27, 47, 28, 9]. Common techniques include warping the first frame of the video and post-processing it with the generative model. In addition, many methods include depth estimation maps to further constrain the generation. All these methods have in common that they use the optical flow mainly as a guiding tool for the V2V or I2V task using sophisticated and complicated mechanisms. In contrast, no other approach allows for generating both unrelated patterns from motion (e.g., waves) and content from prompt (e.g., field of flowers), while being simple.
@@ 3.1 Diffusion Model for Video Generation
Text-to-video generation involves the creation of coherent and high-quality videos conditioned on textual prompts. Due to the high dimensionality and temporal complexity of video data, recent approaches [34] compress the spatial resolution of the input videos $V$ into a lower dimensional latent representation $z$ using an encoder $\mathcal{E}$, i.e., $\mathcal{E}(V)=z$. To project the latent codes back into pixel space, a decoder $\mathcal{D}$ operates on this latent representation to reconstruct the videos.
@@
In this work, we build our architecture on top of AnimateDiff [15]. To effectively capture spatial and temporal dynamics, AnimateDiff first expands Stable Diffusion to video synthesis, inflating the 2D U-net architecture into a 3D U-net. Secondly, it adds temporal convolutions and attention mechanisms to capture temporal dynamics. As a result, the diffusion model factorizes its internal operations over space and time, meaning that each layer operates only on the space or time dimension, not both at the same time. Formally, setting as $B$, $F$, $H\times W$, and $C$ as the batch size, number of frames, spatial dimensions, and channel dimension, respectively, each feature map represents a 5D tensor of $B\times F\times H\times W\times C$ that can be fed into the following layers:
  • spatial layers: each old 2D convolution layer as in the 2D U-net is extended to be space-only 3D convolution over $H\times W\times C$. Each spatial attention block remains as attention operating on individual frames.
  • temporal layers: a temporal attention block is added after each spatial attention block. It performs attention over $F$. The temporal attention block is important to capture good temporal coherence.
@@ 3.2 Optical flow conditioning in T2V Architecture
In this paper, we propose OnlyFlow to modify this traditional T2V scheme by adding motion constraint. Fig. 2 illustrates an overview of our model architecture. More precisely, for enhanced controllability in video generation, we propose to condition the video generation on an optical flow signal $f$. Here, $f$ can be synthetically created or generated through standard optical flow extraction techniques [39].
The architecture of OnlyFlow extends AnimateDiff by incorporating a motion encoder that processes optical flow input and extracts motion features at various scales. We then proceed to inject these signals into each temporal attention layers of the U-net. Let us formally describe both the motion encoder architecture and the feature injection process:
  • Motion feature architecture: inspired by T2I-adapters [30] and CameraCtrl [18], our motion encoder architecture $\Phi$ first unshuffles [7] the input flow $f$ before applying a convolutional layer. It then processes the features into a sequence of spatial resnet blocks and temporal attention operations. Finally, we store the motion features $\Phi(f)=\{c_{k}\}_{k=1}^{4}$ after each temporal operation to inject them into the U-net.
  • Motion feature injection: As seen in Fig. 3, we combine the latent features $h_{k}$ of the $k^{\text{th}}$ block of the U-net with the control signal from motion features $c_{k}$ through element-wise addition. Then, the combined features are processed by a linear layer and scaled using a parameter $\gamma$. Finally, the output is summed back to $h_{k}$ and it is directly fed back into the temporal layers of the U-net architecture. Formally, the complete injection operation for the layer $k$ is:
    $h_{k}^{\prime}=h_{k}+\gamma\;\text{Linear}_{k}\left(c_{k}+h_{k}\right).$ (2)
Here, the parameter $\gamma$ effectively controls the influence of the optical flow from the input auxiliary video onto the generated video. It is the main parameter to boost or decrease the auxiliary motion injection over the generated output. We call $\gamma$ the optical flow conditioning strength.
@@
3.3 Training OnlyFlow
To train our proposed approach, we follow the original formulation of diffusion model training, but include our control signal. More specifically, we optimize the parameters $\phi$ of the motion encoder $\Phi$ by minimizing the standard diffusion model loss
$\mathcal{L}(\phi)=\mathrm{E}_{(V,p),t,\epsilon}\|\epsilon-\epsilon_{\theta}(z_{t},t,p,\Phi(f))\|,$ (3)
where, $z_{0}=\mathcal{E}(V)$, $z_{t}$ is created using Eq. 1, $p$ is the corresponding textual description of $V$, and $f$ is the output of an optical flow predictor $\mathcal{T}$ using as input $V$, i.e., $f=\mathcal{T}(V)$. Of course, during the training phase, we freeze the parameters of AnimateDiff. Only the motion encoder and the merging linear layers in the attention blocks are trained. Additionally, we set the optical flow conditioning strength to $\gamma=1.0$.
4 Experimental Results
4.1 Implementation details
Dataset and Preprocessing
To train OnlyFlow, we used the WebVid dataset [1], which consists of 10.7M video-caption pairs, totaling 52K hours of video content. The majority of these videos have a length between a few seconds and 1 min, with 100k to 500k pixels by frame. All videos are spatially resized to $256\times 384$ pixels, and $F=16$ consecutive frames are randomly chosen, matching the training practice used in AnimatedDiff [15].